Circular No. 6548

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996ce IN MARKARIAN 785 C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of

an apparent supernova (V ∼ 18.0) on a film taken on 1996 Dec. 16.23 UT with the OCA 0.9-m Schmidt telescope. SN 1996ce is located at α = 13^{h}16^{m}15^{s}.47, δ = +30^{o}15′52′′.6 (equinox 2000.0), which is 6′′.0 west and 1′′.9 north of the center of the galaxy Mrk 785 = KUG 1313+305 = MCG +05-31-167; nothing is visible at this location on a film obtained several years ago at OCA. S. Ilovaisky and J. Taupenas (Observatoire de Haute-Provence) confirm the new star with a V CCD observation obtained on 1997 Jan. 26.10 with the 1.2-m telescope. A nearby star (V ∼ 16.0) has position end figures 14^{s}.81, 17′17′′.0.

GRO J1744–28 M. J. Stark, University of Maryland and Goddard Space Flight Center

(GSFC); K. Jahoda and J. H. Swank, GSFC; and T. Strohmayer, GSFC and Universities Space Research Association, report: “We observed GRO J1744–28 on Jan. 18 with the Rossi X-ray Timing Explorer (RXTE) and confirm the earlier report (IAUC 6530) that it is again in outburst. The flux detected by the Proportional Counter Array on RXTE is roughly equivalent to the flux detected from the Crab nebula. The fluxes observed during our first eight days of monitoring are as follows: Jan. 18, 12 800 counts/s; 19, 12 700; 20, 13 100; 21, 13 100; 22, 12 800; 24, 13 500; 25, 13 400; 26, 13 200; 27, 13 500 (where 10 000 counts/s implies a luminosity of 3.5 × 10^{38} erg/s in the range 2–60 keV for this source at 8 kpc). The luminosity during bursts reaches as high as 1.2 × 10^{40} erg/s. The average interval between bursts is 38 min. The burst duration is 10–12 s, which is similar to the duration of bursts on 1996 Mar. 1, when the source was last at this intensity. The current x-ray flux is intermediate between the flux corresponding to the possible infrared detection of a counterpart (x-ray flux ∼ 1.7 Crab) and the flux corresponding to several upper limits (x-ray flux ∼ 0.5 Crab; IAUC 6369; Cole et al. 1997, Ap.J., in press, with preprint avail-

able at http://xxx.lanl.gov/e-print/astro-ph/9610079). Despite the

difficulty of observing near the sun we urge renewed searches for the infrared counterpart.”

1997 January 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green